OpenStudy (friendzone):

(x-1/2)+(x^2+1/3)=1-1/3 result 1, -5/2 x(x-2)+x(x-6)+43=7x-(x+5)(x-5) result 2, 3 x^2-3√3x+5-√3=0 result √3-1, 2√3+1 (x^2+√2-2/2)-(x-√2 / √2)=(x+√2/2) result 0,1+√2

OpenStudy (anonymous):

\frac{x}{x-2}= \frac{2}{x-2}- \frac{2}{3} /*(x-2) multiplying both sides x=2- \frac{2(x-2)}{3} /*3 3x=6-2(x-2) 3x=6-2x+4 /+2x both sides 5x=10 /:5 divide 5 both sides x=2 The result is 2.